OpenStudy (anonymous):

Find sin 2x, cos 2x, and tan 2x from the given information. csc x = 4, tan x < 0

OpenStudy (anonymous):

$$\sin x = \frac 14; \cos x = \sqrt{1- \frac 1{16} }$$ you can find the sin2x and even cos 2x and then tan 2x is easy

OpenStudy (anonymous):

why is it 1/4?

OpenStudy (anonymous):

$$\sin x = \frac1{\csc x} $$

OpenStudy (anonymous):

How did you get these? That's what I'm having trouble with.

OpenStudy (anonymous):

that's the definition of csc x

OpenStudy (anonymous):

trigonometric identities

OpenStudy (anonymous):

it starts with sinx^2+cosx^2=1

OpenStudy (asnaseer):

this might help you: http://www.sosmath.com/trig/Trig5/trig5/trig5.html